OpenStudy (anonymous):

How do I solve for r in w=Cr^-2?

OpenStudy (anonymous):

w = Cr^-2 w = C / r^2 (r^2)(w) = C r^2 = w / C sqrt (r^2) = sqrt (w/C) r = +/- sqrt (w/C)

OpenStudy (anonymous):

Thank you very much!